With what tension must a rope with length 5. 0m and mass 0. 16kg be stretched for transverse waves of frequency 60. 0Hz to have a wavelengths of 0. 80m?
1
Expert's answer
2021-07-25T09:24:25-0400

We know that

"n=\frac{1}{2l}\sqrt{\frac{T}{m}}"

Put value

"60=\frac{1}{2\times5}\sqrt{\frac{T}{0.16}}"

"600=\sqrt{\frac{T}{0.16}}"

T=576N